Subject: APO 3.5-m Users Committee minutes 2/27/06
From: Bruce Gillespie
Submitted: Thu, 2 Mar 2006 12:59:02 -0700
Message number: 951
(previous: 950,
next: 952
up: Index)
APO 3.5-m Users Committee Phonecon, 2/27/06
Attending: Suzanne Hawley, Bruce Gillespie, Michael Strauss, Jon
Holtzman, Remy Indebetouw, Bruce Balick, Russet McMillan, Karl
Glazebrook
Absent: Don York, John Bally, Al Harper
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from Institutional representatives:
Princeton (Michael Strauss)--Michael reported that the PU users are
happy.
Colorado--no report.
NMSU (John Holtzman)--One user reported some positive feedback on how
well her challenging observations of Venus were supported by the site
staff.
Johns Hopkins--no report.
Chicago--no report.
UWashington (Suzanne Hawley)--Suzanne reported that everything is
going fine, plenty of good thesis data being taken. Also, a flare-
star observing run is planned for next month with 5-second time
resolution using DIS.
Univ. of Virginia (Remy Indebetouw)--Remy said he didn't think UVa
was using the telescope very much yet. Suzanne reminded the group
that UVa is getting 5 to 6 half-nights per quarter of DD time. Remy
also wanted to know who to contact about NIC-FPS. BruceG suggested
he contact either Stephane Beland at CU (sbeland@colorado.edu), or
any of the 3.5-m Observing Specialists.
**********************************
Discussion of telescope/instruments report:
BruceG summarized highlights from the report (see below). Suzanne
added that we are planning to have an engineering shutdown in May to
install the new drive boxes.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights,
1/20/06 through 2/22/06
Bruce Gillespie
0) Overview
Highlights include a continued good run of operability for the
telescope and instruments, completion of the new drive-box CIF
project, and progress on the new top-end and DIS-upgrades projects.
The unusually good weather for this time of year persisted, and was
largely well used for observing, although we are beginning to have
concerns about potential fire risks related to the dry conditions.
We are also beginning to plan the summer engineering shutdowns.
1) Telescope
The 3.5-m telescope was largely well-behaved during the past month.
During engineering time, the drive servos were tested and tuned to
try to solve the problem of poor tracking in the parts of the sky
where azimuth velocity approached zero and reverses (in the NW and NE
quadrants). Some improvement was noted, but it is felt that fixing
the problem will only be possible when we upgrade and replace the
drive servos, hopefully next year.
2) The performance of the instruments was near nominal, and
noteworthy are:
o DIS: Testing of wavelength stability with instrument rotation
was done, which indicated a roughly 1-Angstrom drift in the spectrum
at various rotator angles. Discussion and analysis of this is ongoing.
o Echelle: The echelle instrument control computer was upgraded
without incident.
o NIC-FPS: Last-minute problem with dewar vacuum, currently
under investigation.
3) Engineering and CIF projects
The assembly of the three new drive boxes was completed, and they
passed their bench testing. The boxes will be installed on the
telescope during the summer, or earlier if there is a problem with
the existing boxes. Design work continued on the new top end, and
materials have been received and machining work on some of the parts
continues. Maintenance and enhancement of TUI and other software
systems continued. The DIS upgrade project (gratings, optics, new
detector) is underway; parts are being ordered.
4) Miscellaneous
Gabrelle Saurage (the new Observing Specialist) spent a training
weekend at the telescope.
**********************************
Short-term data storage policy:
For the record, Suzanne summarized our previous discussion and
decisions on 3.5-m data storage. APO will not archive 3.5-m data.
As has always been our practice, telescope users are obliged to
download their data to their home institutions as soon as is
practical. We do intend to keep the data from the past 6 to 9 months
on disk at the site. At the start of every quarter, the data older
than 6 months will be erased. If we end up short on disk space, this
policy may be revised as necessary.
There were no further comments or discussion on this topic.
**********************************
GRB alert program status:
Suzanne mentioned that the ARC-wide GRB alert program has been
suspended by mutual agreement with the GRB team. It is still
possible for the individual ARC institutions to make GRB follow-up
observations within their internal scheduled time allocations.
**********************************
Backup program advisement:
BruceG offered the committee a draft paragraph to be added to the
automated observer notification message. After some discussion about
the wording and intent, BruceG modified the paragraph as follows. It
was also suggested that the advisement be added to the proposal form.
Draft text: "There may be occasions where unanticipated observing
conditions or hardware/software problems prevent you from making your
primary program observations. We advise all 3.5-meter telescope
users that we strongly recommend that you have in reserve at least
one backup science observing program when you use the telescope,
preferably one that uses a different instrument from your primary
program. If you do not have a backup program and you determine that
your primary observing program is not achievable, you should consider
relinquishing the telescope to the observatory staff who will attempt
to make the best alternative use of the observing time that
conditions and circumstances allow."
**********************************
Triplespec white paper:
Suzanne noted that UVa has fulfilled an ARC request to provide a
short white paper describing the Triplespec NIR spectrograph project,
written by John Wilson. The white paper was distributed to the Users
Committee members prior to the meeting, and Suzanne asked the
committee members to distribute it to their respective user
communities for comments. The committee members are to forward any
feedback to Suzanne within a week, from which she will report our
assessment to the ARC Board of Governors.
**********************************
Actions from previous meetings:
ACTION: Once we collect the the names of the designated
institutional systems administrators, BruceG will have them listed on
the APO web pages for reference. STATUS: Completed, see http://
www.apo.nmsu.edu/Site/directory/projdirectory/default.htm.
ACTION: BruceG will write a paragraph about the importance of having
back-up programs, to be added to the e-mail automated message we send
observers prior to their runs. STATUS: In draft.
ACTION: Users Committee members should poll their users to ask if
the the present data distribution/security systems at APO are
adequate, and if not, why. STATUS: Completed.
ACTION: Users Committee members should poll their users to ask if we
should permanently archive all 3.5-m data, and possibly make it
available to the public, and if so, why. STATUS: Completed.
[new actions from last month]:
ACTION: The Users Committee members will need to help disseminate
the [Futures Committee] draft report to their users. STATUS: Report
not yet available.
ACTION: BruceG will ask Russell what the pros and cons are of
keeping TUI compatible with older versions of operating systems and
libraries. STATUS: Discussed with RO.
ACTION: Users Committee members should poll their users to ask if
there are compelling reasons to keep TUI operable without using the
latest version of operating systems. STATUS: Open. We are
considering a policy where TUI users will need to work remotely from
computers that have at least the "n-minus-one" versions of operating
systems and libraries.
**********************************
Next Users Committee phonecon will be on Tuesday, 4 April, at 8:30 AM
Pacific Time. Agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 951 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO